# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

### v1

@@ H.5 Ablations for VERSE (Step 2)
gpt-4o 0.49 0.53 0.63
w/ CoT 0.46 0.50 0.67
claude-3-5-sonnet 0.38 0.41 0.63
w/ CoT 0.51 0.54 0.71
claude-3-5-haiku 0.38 0.41 0.75
w/ CoT 0.39 0.42 0.66
gpt-4o-mini 0.37 0.40 0.63
w/ CoT 0.35 0.37 0.63
claude-3-haiku 0.20 0.21 0.81
w/ CoT 0.19 0.20 0.87
gemma-2-27b 0.18 0.19 0.70
w/ CoT 0.15 0.16 0.76
gemma-2-9b 0.09 0.09 0.73
w/ CoT 0.18 0.19 0.88
llama-3-8b 0.05 0.06 1.06
w/ CoT 0.15 0.16 0.92
Table 12: (Step 2) Performance of different models with and without Chain-of-Thought (CoT) prompting. Better results for each model pair are shown in bold. The effect of CoT is inconsistent and depends on the specific models. Some models do benefit from CoT.
@@
As seen in Table 12, chain of thought reasoning doesn’t GPT series models much, in fact, it hurts. However, for smaller models, doing chain of thought boosts the performance from completely random to 0.20.2~{}0.20.2. However, even for the models that do benefit from CoT (in terms of correlation), some of them end up having higher MSE, which means it become biased towards giving higher scores to machine translations.
Default No Crit No Human Tr
@@
gpt-4o 0.49 0.53 0.63 0.07 0.08 1.22 0.51 0.56 0.60
gpt-4o-mini 0.37 0.40 0.63 0.03 0.04 1.02 0.40 0.44 0.62
gemma-2-27b 0.18 0.19 0.70 0.10 0.11 0.70 0.24 0.25 0.75
gemma-2-9b 0.09 0.09 0.73 0.00 0.00 0.83 0.16 0.17 0.76
llama-3-8b 0.05 0.06 1.06 -0.02 -0.03 0.73 0.05 0.05 1.06
Table 13: (Step 2) Performance comparison across different models and configurations. No Crit refers to when human written rubric was not provided and No Human Tr means human translation was not provided. Notice that while having access to human translation is not critical for large models, having access to human-curated rubic is essential for all models.
@@
20 0.49 0.53 0.63
gpt-4o-mini 0 0.37 0.40 0.63
5 0.21 0.23 0.75
10 0.15 0.16 0.75
15 0.23 0.25 0.72
20 0.20 0.22 0.78
llama-3-8b-it 0 0.05 0.06 1.06
5 0.05 0.05 1.53
10 0.09 0.10 1.27
15 0.06 0.07 0.82
20 0.06 0.06 0.77
Table 14: (Step 2) Performance comparison across k shot examples provided (in no COT setting) in answering LLM generated evaluation questions. The best results for each model are shown in bold. Note that gemma models are excluded because of their shorter context length.
@@
As Table 14 shows, providing a few shot examples doesn’t significantly boost performance across models and for one model (gpt-4o-mini) there is a decline in model performance.
@@ I.1 Translation Generation
Figure 8: Fine-grained evaluation of 8 stories using VERSE. Notice that models score widely differently for different stories. The difference is more pronounced for less powerful model, for whom GPT4o $n$ refers to the number of questions belonging to that category. The legend on the bottom left shows which color belongs to which model.
@@ I.2 LLM’s tendency to overestimate scores
As shown in Table 15 and 16, evaluator LLM GPT4-o tends to overestimate the quality of LLM generated translations over human translations. An interesting pattern is observed with the direction of overestimation. For DA-MQM (Step 1) (Table 15) scores, providing reference translation decreases the average score, while for LLM generated QnA in Step2 (Table 16), the average score increases. This is in line with the results observed in the ablations in Table 10 and 13, where MSE increases when reference is not provided for DA-MQM (Step 1), and decreases for VERSE (Step 2).
@@
Notice, however, that the relative ranks of the models do change in a few cases in these two settings, especially in the middle ranks. The best and worst-performing models’ ranks do not change.
This observation supports our previous observation in Section 6 that GPT4-o as a judge can differentiate between obviously good and obviously bad translations, however, its capability to detect differences in similarly (high) quality translations is not as robust.

### v2

@@ H.5 Ablations for VERSE (Step 2)
gpt-4o 0.49 0.54 0.61
w/ CoT 0.45 0.50 0.69
claude-3-5-sonnet 0.51 0.54 0.73
w/ CoT 0.36 0.39 0.65
claude-3-5-haiku 0.37 0.40 0.77
w/ CoT 0.39 0.42 0.67
gpt-4o-mini 0.37 0.40 0.61
w/ CoT 0.34 0.36 0.63
claude-3-haiku 0.20 0.21 0.82
w/ CoT 0.17 0.19 0.89
gemma-2-27b 0.13 0.14 0.71
w/ CoT 0.09 0.10 0.78
gemma-2-9b 0.09 0.09 0.71
w/ CoT 0.16 0.17 0.88
llama-3-8b 0.05 0.05 1.08
w/ CoT 0.14 0.16 0.90
Table 12: (Step 2) Performance of different models with and without Chain-of-Thought (CoT) prompting. The effect of CoT is inconsistent and depends on the specific models. Some models do benefit from CoT.
@@
As seen in Table 12, chain of thought reasoning doesn’t GPT series models much, in fact, it hurts. However, for smaller models, doing chain of thought boosts the performance from completely random to 0.20.2~{}0.20.2. However, even for the models that benefit from CoT (in terms of correlation), some of them have higher MSE, which usually means they tend to give higher scores to machine translations.
Default No Rubric No Human Tr
@@
gpt-4o 0.49 0.54 0.61 0.08 0.09 1.22 0.52 0.57 0.59
gpt-4o-mini 0.37 0.40 0.61 0.02 0.02 1.05 0.42 0.46 0.60
gemma-2-27b 0.13 0.14 0.71 0.07 0.08 0.70 0.25 0.26 0.75
gemma-2-9b 0.09 0.09 0.71 0.02 0.02 0.79 0.17 0.18 0.75
llama-3-8b 0.05 0.05 1.08 -0.02 -0.02 0.73 0.06 0.06 1.08
Table 13: (Step 2) Performance comparison across different models and configurations. No Rubric refers to when human written rubric was not provided and No Human Tr means human reference translation was not provided. Notice that while having access to human translation is not critical for large models, having access to human-curated rubic is essential for all models.
@@
20 0.50 0.54 0.63
gpt-4o-mini 0 0.37 0.40 0.61
5 0.20 0.21 0.75
10 0.14 0.15 0.74
15 0.20 0.22 0.73
20 0.19 0.21 0.77
llama-3-8b-it 0 0.05 0.05 1.08
5 0.04 0.04 1.59
10 0.06 0.07 1.31
15 0.06 0.06 0.82
20 0.05 0.05 0.77
Table 14: (Step 2) Performance comparison across k shot examples provided (in no COT setting) in answering LLM generated evaluation questions. Note that gemma models are excluded because of their shorter context length.
@@
As Table 14 shows, providing a few shot examples doesn’t significantly boost performance across models.
@@ I.1 Translation Generation
Figure 8: Fine-grained evaluation of 8 stories using VERSE. Notice that models score widely differently for different stories. The difference is more pronounced for less powerful model, for whom evaluator LLM is more sensitive. $n$ refers to the number of questions belonging to that category. The legend on the bottom left shows which color belongs to which model.
@@ I.2 LLM’s tendency to overestimate scores
As shown in Table 15 and 16, evaluator LLM GPT4-o tends to overestimate the quality of LLM generated translations over human translations. An interesting pattern is observed with the direction of overestimation. For RULER (Step 1) (Table 15) scores, providing reference translation decreases the average score, while for LLM generated QnA in Step2 (Table 16), the average score increases. This is in line with the results observed in the ablations in Table 10 and 13, where MSE increases when reference is not provided for RULER (Step 1), and decreases for VERSE (Step 2).
@@
Notice, however, that the relative rank of the models changes in a few cases in these two settings, especially in the middle ranks. However, the best and worst-performing models’ ranks do not change whether reference translation is provided or not.
This observation supports our previous claim in Section 6 that GPT4-o as a judge can differentiate between obviously good and obviously bad translations, however, its capability to detect differences in similarly (high) quality translations is not as robust.